Rachael runs 2 kilometers to a bus stop, and then rides 4.5 km to school. On average the bus is 45 km/hr faster than Rachael's average running speed. If the entire trip takes 25 minutes, how fast does Rachael run?
Answer by stanbon(75887)

Rachael runs 2 kilometers to a bus stop, and then rides 4.5 km to school. On average the bus is 45 km/hr faster than Rachael's average running speed. If the entire trip takes 25 minutes, how fast does Rachael run?
-------
Running DATA:
distance: 2 km ; rate = x km/h ; time = 2/x hrs
-------
Bus DATA:
distance: 4.5 km ; rate = x+45 km/h ; time = 4.5/(x+45) hrs.
----------------------
Equation:
time + time = 25/60 hr
----
2/x + 4.5/(x+45) = 5/12
----
Multiply thru by 12x(x+45) to get:
24(x+45) + 12*4.5x = 5x(x+45)
---------------
24x + 1080 + 54x = 5x^2 + 225x
=============================
78x + 1080 = 5x^2 + 225x
5x^2 + 147x - 1080 = 0
-----------------------
Positive solution:
x = 6.0868 km/h (running rate)
=================
Cheers,
Stan H.